which trinomial is equivalent to (3x - 2) (x +4)


1
Expert's answer
2022-05-11T15:57:48-0400

We'll use the FOIL method.

The FOIL method is a way of systematically multiplyilng a binomial with another binomial. This is done by multiplying the First terms with each other, then the Outer terms of the binomials, the Inner terms and then the Last terms.

Then we sum up the partial products of the two binomials beginning from the first, outer, inner, and then the last.

"(3x-2) (x+4) =\\
=3x\cdot x+3x\cdot 4+(-2)\cdot x+(-2)\cdot 4=\\
=3x^2 +12x-2x-8=\\
=3x^2 +10x-8."